### v1

# Can LLMs Alleviate Catastrophic Forgetting in Graph Continual Learning? A Systematic Study

Abstract
Nowadays, real-world data, including graph-structure data, often arrives in a streaming manner, which means that learning systems need to continuously acquire new knowledge without forgetting previously learned information. Although substantial existing works attempt to address catastrophic forgetting in graph machine learning, they are all based on training from scratch with streaming data. With the rise of pretrained models, an increasing number of studies have leveraged their strong generalization ability for continual learning. Therefore, in this work, we attempt to answer whether large language models (LLMs) can mitigate catastrophic forgetting in Graph Continual Learning (GCL). We first point out that current experimental setups for GCL have significant flaws, as the evaluation stage may lead to task ID leakage. Then, we evaluate the performance of LLMs in more realistic scenarios and find that even minor modifications can lead to outstanding results. Finally, based on extensive experiments, we propose a simple-yet-effective method, Simple Graph Continual Learning (SimGCL), that surpasses the previous state-of-the-art GNN-based baseline by around 20% under the rehearsal-free constraint. To facilitate reproducibility, we have developed an easy-to-use benchmark LLM4GCL for training and evaluating existing GCL methods. The code is available at: https://github.com/ZhixunLEE/LLM4GCL.
1 Introduction
Graph Machine Learning (GML) has become a cornerstone of modern data science, enabling intelligent systems to analyze complex relational structures across domains such as social networks fan2019graph , financial networks han2025mitigating ; li2022devil ; yang2025grad , recommendation systems li2025g ; wu2022graph , and biological structures.chen2023uncovering ; wang2022molecular . Currently, with the remarkable capability to model diverse graph structures and node attributes, Graph Neural Networks (GNNs) have emerged as a prevalent approach for GML tasks. However, most existing approaches assume a static setting where the entire dataset and label space are known in advance. In contrast, real-world graphs are dynamic in nature, as new nodes, edges, and even previously unseen classes may appear over time. This evolving structure introduces the need for Graph Continual Learning (GCL), where models must incrementally integrate new knowledge while retaining previously learned information. Although significant progress has been made in GCL, with numerous studies liu2021overcoming ; liu2023cat ; niu2024graph ; zhang2023continual ; zhang2022hierarchical ; zhang2022sparsified ; zhang2023ricci ; zhou2021overcoming , substantial challenges still remain unresolved:
  • A Lack of investigation into the rationality of experimental setups in GCL. This paper focuses specifically on the more mainstream and challenging setting of Node-level Class-Incremental Learning (NCIL). Based on the scope of each task’s test set, we divide NCIL into two categories: local testing and global testing. Although numerous works have been proposed within each category, there is currently no study that analyzes the rationality of the setups.
  • Existing methods are all trained from scratch on streaming graph-structure data. The great success of pretrained models like GPT achiam2023gpt and CLIP radford2021learning in fields such as natural language processing and computer vision has provided new opportunities for continual learning. However, due to a lack of powerful foundation models in GML, existing baselines are all trained from scratch.
  • No existing studies that evaluate the performance of LLM-based methods in GCL. Currently, LLM-based methods have achieved state-of-the-art performance in node classification tasks li2024glbench ; wu2025comprehensive , significantly outperforming traditional GNN models. However, there is a lack of systematic benchmark evaluations assessing the performance of these models in the GCL setting.
To fill these gaps, we systematically analyze the NCIL task and explore the opportunities for mitigating catastrophic forgetting in GCL via LLMs. Specifically, we first identify a flaw in the local testing setting of NCIL, where task ID leakage occurs. This causes class-incremental learning to degrade into task-incremental learning, significantly reducing the task’s difficulty. We employ an extremely simple method to predict the task ID with 100% accuracy, achieving a forget-free performance and matching the results of the previous SOTA model niu2024replay . Subsequently, we systematically evaluate the performance of LLMs and Graph-enhanced LLMs (GLMs) in global testing, a more realistic and challenging setting. We find that by incorporating common techniques such as Parameter Efficient Fine-Tuning (PEFT) or prototype classifiers, it is possible to leverage the strong generalization ability of pretrained models to surpass existing baseline methods. Finally, based on extensive experiments, we propose a simple yet effective method called SimGCL (Simple Graph Continual Learning for short). Technically, to better incorporate graph structural knowledge, we design ego-graph-derived prompts for each node, which effectively capture textual and structural features of the nodes. Then, in the first session, we applied LoRA hu2022lora to efficiently fine-tune the LLMs with instruction tuning, aiming to narrow the distribution gap between the specific dataset and the pretrained dataset. In the incremental sessions, we utilize a training-free prototype classifier for prediction, avoiding parameter updates caused by the arrival of new knowledge, which greatly alleviates catastrophic forgetting. Based on experiments in NCIL and Few-Shot Node-level Class-Incremental Learning (FSNCIL), our proposed SimGCL can significantly surpass the current SOTA models.
In summary, our main contributions are as follows:
  • Analysis of the rationality of experimental setups. This paper is the first to analyze the flaws in certain experimental setups in GCL and prevent task ID leakage.
  • Comprehensive benchmark. We propose LLM4GCL, the first comprehensive benchmark for LLMs on GCL. We integrate 9 LLM-based methods and 7 textual-attributed graphs for evaluations.
  • Model design. We propose a simple yet effective model, SimGCL, which is able to surpass all existing baselines and achieve an absolute increase of nearly 20% on certain datasets.
  • Easy-to-use platform. To facilitate future work and help researchers quickly use the latest models, we develop an easy-to-use open-source platform. We hope to encourage more researchers to explore the use of pretrained models to mitigate catastrophic forgetting in GCL. The code is available at: https://github.com/ZhixunLEE/LLM4GCL.
2 Formulations and Background
Text-attributed Graphs (TAGs). In this benchmark, we evaluate the continual learning capabilities of LLMs through node classification tasks on Text-Attributed Graphs (TAGs) ma2021deep . Formally, a TAG is represented as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, where $\mathcal{V}$ is the set of $N$ nodes, $\mathcal{E}$ is the edge set, and $\mathcal{R}$ is the collection of textual descriptions associated with each node $v_{i}\in\mathcal{V}$. Each description $r_{v_{i}}\in\mathcal{D}^{L_{i}}$ consists of tokens from a vocabulary $\mathcal{D}$ with sequence length $L_{i}$. Traditionally, the textual attributes of nodes can be encoded into shallow embeddings as $\textbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}]\in\mathbb{R}^{N\times d}$ using techniques like bag-of-words or TF-IDF salton1988term , where $d$ is the embedding dimension. Such transformation is widely adopted in the graph neural network literature. In contrast, LLM-based approaches directly process the raw text. Given a TAG with labeled nodes $\mathcal{V}^{l}$, the node classification task aims to predict the labels of the unlabeled nodes $\mathcal{V}^{u}$.
Node-Level Class-Incremental Learning (NCIL). Node-Level Class-Incremental Learning (NCIL) represents a fundamental paradigm in GCL, addressing incremental node classification tasks in an expanding graph $\mathcal{G}$. This paradigm well reflects the real-world applications, where new categories (e.g., new products for recommendation systems or novel interdisciplinary techniques for citation networks) are continually incorporated into the graph in a streaming manner as the domain evolves. As shown in Figure 1, the ordered training task sequence $\{\mathcal{T}_{s_{1}},...,\mathcal{T}_{s_{n}}\}$ consists of tasks $\mathcal{T}_{s_{i}}=\{\mathcal{G}_{s_{i}},\mathcal{C}_{s_{i}},\mathcal{V}_{s_{i}}^{l}\}$, where each $\mathcal{G}_{s_{i}}\subset\mathcal{G}$ represents the task subgraph, $\mathcal{C}_{s_{i}}$ denotes the class set, and $\mathcal{V}_{s_{i}}^{l}$ contains the labeled nodes, with $\mathcal{G}_{s_{i}}\cap\mathcal{G}_{s_{j}}=\mathcal{C}_{s_{i}}\cap\mathcal{C}_{s_{j}}=\varnothing$ for all $i\neq j$. During training, the model is expected to undergo incremental training on each $\mathcal{T}_{s_{i}}$ and make predictions of the unlabeled nodes.
Few-Shot Node-Level Class-Incremental Learning (FSNCIL). Newly emerged categories always contain a few samples. Consequently, it is difficult to collect sufficient samples for incremental tasks that match the initial task’s scale, leading to the FSNCIL paradigm. Similar to NCIL, the FSNCIL paradigm also operates on a task sequence $\{\mathcal{T}_{s_{0}},\mathcal{T}_{s_{1}},...,\mathcal{T}_{s_{n-1}}\}$, where $\mathcal{T}_{s_{0}}$ serves as the base session and $\mathcal{T}_{s_{i}}$ ($i\in\{1,2,\ldots,n-1\}$) are incremental sessions. Moreover, incremental sessions in FSNCIL adhere to the $m$-way $k$-shot setting, where the number of classes $m$ and the sample size of each class $k$ are significantly smaller than in the base session. During training, the model is first trained on the base session, then fine-tuned on the incremental session sequence. It is noteworthy that FSNCIL is more challenging compared to NCIL, as the limited sample size exacerbates the risk of overfitting.
Figure 1: Left: The settings of LLM4GCL. During training, the model receives only the current task’s graph, while testing employs the complete graph. Furthermore, GFSCIL is restricted to a few-shot node samples during its training phase. Right: The evaluation baselines of LLM4GCL, which can be categorized into three types: GNN-based, LLM-based, and GLM-based models.
3 LLM4GCL: A Systematic Study of LLMs for Graph Continual Learning
3.1 Learning Scenarios
LLM4GCL supports both NCIL and FSNCIL paradigms. Prior to this benchmark, existing GCL methods prevalently adopted the evaluation metrics from CGLB zhang2022cglb , which we refer to as the local testing. Specifically, when the $i$-the task $\mathcal{T}_{s_{i}}$ has been trained and the the model need to be evaluated, this testing manner has an ordered testing task sequence $\{\mathcal{T}_{q_{1}},\ldots,\mathcal{T}_{q_{i}}\}$, where $i\leq n$. The evaluation process consists of $i$ independent evaluations, with each testing task $\mathcal{T}_{q_{j}}=\{\mathcal{G}_{q_{j}},\mathcal{C}_{q_{j}}\}$ ($j\leq i$) leveraging the local subgraph ($\mathcal{G}_{q_{j}}=\mathcal{G}_{s_{j}}$) and global class set ($\mathcal{C}_{q_{j}}=\bigcup_{z=1}^{j}\mathcal{C}_{s_{z}}$), respectively. However, we highlight that this setup has a fundamental flaw: for each task, since testing and training samples are drawn from the same subgraph, their distribution similarity makes it trivial to accurately predict task ID, degrading class-incremental to task-incremental learning. For instance, TPP niu2024replay transductively captures task-specific prototypes utilizing a Laplacian smoothing-based matching approach, achieving 100% task ID prediction accuracy and 0% forgetting ratio. However, as demonstrated in Figure 2, our analysis reveals that even using a basic mean pooling operation to get prototypes achieves flawless task ID prediction. Furthermore, with accurate task identification, the model can also attain substantial performance by discarding the GNN entirely and simply employing task-specific two-layer MLPs for classification [^1] For comprehensive details of TPP and experimental setup, please see Appendix A.1.. Consequently, we argue that this evaluation method is fundamentally flawed, as it fails to accurately assess the model’s genuine continual learning capability. In this paper, we primarily investigate the more challenging and realistic global testing setup. In global testing, we also have an ordered testing task sequence, but the difference is that for each testing task $\mathcal{T}_{q_{j}}=\{\mathcal{G}_{q_{j}},\mathcal{C}_{q_{j}}\}$ ($j\leq i$), the evaluated graph is the union of the subgraphs from all previous tasks ($\mathcal{G}_{q_{j}}=\bigcup_{z=1}^{j}\mathcal{G}_{s_{z}}$) and the class set is still global ($\mathcal{C}_{q_{j}}=\bigcup_{z=1}^{j}\mathcal{C}_{s_{z}}$). This setup not only better reflects real-world scenarios but also prevents the issue of task ID leakage, enabling authentic assessment of continual learning performance.
In addition to evaluation settings, previous knowledge leakage and label imbalance also constitute significant concerns for both NCIL and FSNCIL approaches. Since GNNs’ message-passing mechanisms inherently aggregate neighborhood information, these edges connecting old and new task nodes potentially expose previous task knowledge during new task training zhang2022cglb . Given that real-world scenarios often prohibit access to previous task data due to privacy or storage constraints, our benchmark excludes inter-task edges, using only intra-task connections. As for the label imbalance issue, since each session contains only a limited number of classes, while sample sizes vary substantially across classes, the imbalance issue becomes pronounced in continual learning settings. To prevent performance degradation caused by this issue, our benchmark removes the classes with insufficient samples and utilizes a unified sample size for each remaining class.
Figure 2: Left: Task-averaged accuracy comparison of TPP (Laplacian-smoothed prototypes), Mean Pooling (node feature averaging), and MLP (pure MLP predictions). Right: Testing paradigms of local testing (sequential evaluation) versus global testing (whole-graph evaluation).
3.2 Benchmark Settings
Datasets. To ensure comprehensive evaluation, we gather seven diverse and representative datasets: Cora he2023harnessing , Citeseer chen2024exploring , WikiCS liu2023one , Photo yan2023comprehensive , Products hu2020open , Arxiv-23 he2023harnessing , and Arxiv hu2020open , according to the following criteria: (1) Multiple domains. LLM4GCL incorporates datasets spanning multiple domains, including citation networks, web link networks, and e-commerce networks, reflecting diverse real-world applications; (2) Diverse scale and density. LLM4GCL datasets cover a wide range of scales, from thousands to hundreds of thousands of nodes. Moreover, the density of these datasets also varies significantly, supporting a robust evaluation of the model’s effectiveness; (3) Various sessions. LLM4GCL datasets provide varied session settings of both total session number and classes per session, allowing comprehensive analysis of the model’s continual learning performance. The statistics of these datasets and descriptions are provided in Appendix A.2.
[…]
Metrics. In LLM4GCL, we employ Accuracy as the evaluation metric for each node classification task. Following established continual learning benchmark rebuffi2017learning , we define $\mathcal{A}_{i}$ as the accuracy after the $i$-th training task. The final metrics for both NCIL and FSNCIL are $\bar{\mathcal{A}}=\frac{1}{n}\sum_{i=1}^{n}\mathcal{A}_{i}$ and $\mathcal{A}_{N}$, which refer to the average accuracy across all tasks and the accuracy after the last task, respectively. These two metrics can well assess the model’s overall performance and ultimate knowledge preservation.
Figure 3: Our proposed SimGCL framework executes two stages: (1) fine-tuning with ego-graph-derived prompts and instructions, followed by (2) training-free prototype generation and cosine similarity-based classification using LLM-generated node embeddings.
3.3 SimGCL: A Simple yet Effective GLM-based Approach for Graph Continual Learning
Currently, the application of LLMs and GLMs to GCL tasks remains nascent. To the best of our knowledge, there are no existing LLM- or GLM-based methods specifically designed for GCL tasks. While directly employing GLMs offers a straightforward solution by combining evolving graph structures with LLMs’ generalization capabilities, this approach incurs substantial computational overhead. Furthermore, as new tasks accumulate, LLMs’ general knowledge progressively degrades, resulting in a biased preference toward newer tasks. Consequently, we propose SimGCL, as depicted in Figure 3, which demonstrates two advantages: (1) Efficiency. SimGCL only requires a single round of instruction tuning in the first session, dramatically reducing training costs. (2) Generalization. SimGCL constructs class prototypes directly from the embeddings produced by its fine-tuned GLM. This training-free approach avoids parameter updates, thereby mitigating catastrophic forgetting.
Instruction Tuning. The instruction tuning phase facilitates LLM adaptation to graph-structured tasks by bridging the distributional gap between the pretraining corpus and target task datasets. Inspired by wang2025exploring , we employ a neighborhood-aware graph prompt to guide the LLM’s processing. Derived from each node’s ego-graph $\tilde{\mathcal{G}}_{v}$, this prompt template comprises three components: (1) Task description outlining the scenario and dataset; (2) Graph structure incorporating target node attributes and neighborhood structure, with each node represented in the format of [Node ID][Node Text]; and (3) Question that specifies requirements and label text options. Moreover, to ensure training efficiency, we limit LLM tuning to the first session and adopt LoRA hu2022lora instead of full-parameter fine-tuning to prevent overfitting. The training objective is to maximize the correct answer probability:
$\max\limits_{\theta}p(y\mid r_{\text{prompt}},\tilde{\mathcal{G}_{v}})=\prod_{l=1}^{L}p(y_{l}\mid y_{<l},r_{\text{prompt}},r_{v},\mathcal{R}_{\mathcal{N}(v)})$ (1)
where $\theta$ is the LoRA parameters, $r_{\text{prompt}}$, $r_{v}$, and $\mathcal{R}_{\mathcal{N}(v)}$ denote the prompt template, target node text, and neighbor text set respectively, with $y_{l}$ being the $l$-th label token and $y_{<l}$ its preceding tokens.
Prototype Generation and Classification. Following instruction tuning, SimGCL progressively generates class prototypes across sessions. Formally, let $f_{\theta}:(r_{\text{prompt}},\mathcal{G}_{v})\mapsto\mathbf{h}_{v}$ denote the LLM’s mapping from inputs to final-layer embeddings, the prototype $\mathbf{c}_{i}$ for class $i$ is computed as:
$\mathbf{c}_{i}=\frac{1}{K}\sum_{j=1}^{|\mathcal{Y}_{b}|}\mathbb{I}(y_{j}=i)\mathbf{h}_{j}$ (2)
where $K=\sum_{j=1}^{|\mathcal{Y}_{b}|}\mathbb{I}(y_{j}=i)$, with $|\mathcal{Y}_{b}|$ being the number of labeled nodes in session $b$ where class $i$ appears, and $\mathbb{I}$ denotes the indicator function. Then during prediction, for node $v_{i}$ with its embedding $\mathbf{h}_{i}$, its probability of belonging to class $j$ is formulated as follows:
$\hat{y}_{i,j}=\frac{\mathbf{h}_{i}\cdot{\mathbf{c}_{j}}}{\left\|\mathbf{h}_{i}\right\|\cdot\left\|\mathbf{c}_{j}\right\|}\cdot\tau$ (3)
where $\tau(\tau>0)$ is the scaling hyperparameter controlling the weight distribution. This training-free approach effectively preserves class-discriminative features across sessions, simultaneously reducing computational cost for new sessions while alleviating catastrophic forgetting in LLMs.
4 Experiments and Analysis
In this section, we conduct extensive experiments to analyze the performance of various GNN-based, LLM-based, and GLM-based methods on GCL tasks. Specifically, Table 1 and Table 2 present the experimental results of these methods under the NCIL and FSNCIL scenarios, respectively; Figure 4 illustrates the performance of SimpleCIL and SimGCL when utilizing LLM models of varying sizes; Figure 3 presents the performance trends of these baseline methods across different session numbers in the Arxiv dataset. Based on these results, we give the following key observations (Obs.):
Table 1: Performance comparison of GNN-, LLM-, and GLM-based methods in the NCIL scenario of LLM4GCL. The best, second, and third results are highlighted.
Methods Cora Citeseer WikiCS Photo Products Arxiv-23 Arxiv Avg.
$\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ Rank
GCN 57.0 38.2 52.4 30.2 54.9 34.9 46.5 19.9 25.5 5.4 19.9 2.9 21.2 1.5 9.7
EWC 56.0 31.0 45.9 28.5 55.3 33.0 46.9 20.5 29.4 15.9 25.7 15.0 24.9 13.8 8.1
LwF 55.7 30.8 47.8 28.2 55.0 34.0 45.8 20.4 26.0 7.5 21.1 6.8 18.9 1.6 10.4
Cosine 65.4 45.2 50.7 31.0 66.5 53.5 63.6 49.6 36.1 16.1 36.1 24.2 37.4 27.8 3.8
TPP 45.7 13.7 45.4 9.6 35.1 9.8 36.3 5.7 15.0 0.0 12.2 0.3 19.6 3.4 14.4
BERT 56.0 29.9 53.8 28.7 58.4 30.0 43.4 18.4 26.9 4.1 27.1 5.1 26.0 9.7 9.1
RoBERTa 54.6 29.6 54.1 28.6 55.1 30.8 43.5 19.1 27.6 3.2 23.1 0.8 24.4 1.4 10.9
LLaMA 65.6 53.8 55.7 31.7 55.5 30.9 44.6 19.1 29.8 0.4 24.3 1.0 24.9 1.4 7.8
SimpleCIL 70.8 58.3 66.4 49.5 71.4 57.3 62.1 52.5 66.8 52.6 52.4 38.8 50.6 36.5 1.9
[…]
ENGINE 59.2 31.3 53.5 29.8 56.4 30.1 47.9 21.0 27.2 1.1 22.5 0.7 24.6 1.3 9.4
GraphPrompter 61.9 46.8 60.2 30.6 59.6 38.3 51.5 31.0 29.0 0.8 23.4 0.9 24.8 1.4 6.7
GraphGPT 55.5 31.6 60.0 30.1 62.0 49.2 50.8 30.2 35.5 3.2 30.7 5.8 32.2 3.9 5.9
LLaGA 58.2 30.2 51.3 27.8 53.7 27.6 47.2 20.7 25.7 0.2 26.9 4.1 26.1 1.3 10.5
SimGCL (Ours) 84.6 80.0 77.1 66.3 73.5 61.9 82.1 72.6 71.1 60.2 38.7 13.6 59.9 33.8 1.4
Table 2: Performance comparison of GNN-, LLM-, and GLM-based methods in the FSNCIL scenario of LLM4GCL. The best, second, and third results are highlighted.
Methods Cora Citeseer WikiCS Photo Products Arxiv-23 Arxiv Avg.
$\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ Rank
GCN 68.0 38.1 39.5 17.4 62.4 47.4 58.5 32.3 36.0 14.1 22.2 5.4 25.3 2.0 9.7
EWC 59.0 36.3 49.2 21.2 58.4 40.4 62.0 28.4 45.7 31.5 36.4 29.5 31.5 18.3 8.3
LwF 63.3 43.5 45.1 20.7 59.4 41.0 60.1 29.4 50.3 38.7 30.8 15.1 27.8 3.8 7.9
Cosine 72.6 57.8 49.1 25.7 68.0 50.7 67.9 50.5 50.9 33.6 40.1 27.9 27.2 17.2 4.8
TEEN 60.9 40.3 59.0 39.5 59.3 42.4 59.3 35.5 49.6 28.4 39.2 27.1 31.8 18.6 6.3
TPP 39.0 9.1 37.3 12.6 37.3 12.4 40.0 14.0 14.0 0.0 7.3 0.0 11.1 6.0 15.3
BERT 56.4 34.7 61.1 38.8 61.7 33.0 47.7 25.8 22.4 3.8 15.2 3.4 14.9 6.0 11.3
RoBERTa 59.6 41.9 54.0 29.2 67.2 42.3 58.2 29.6 38.8 6.9 22.6 1.4 25.7 1.3 9.6
LLaMA 72.6 55.6 75.9 55.5 65.2 43.9 61.4 32.3 43.5 13.7 23.5 1.5 22.7 1.4 7.1
SimpleCIL 69.6 53.6 64.1 49.9 73.2 63.1 66.3 53.0 65.6 53.6 49.8 40.0 46.4 36.6 2.4
[…]
ENGINE 52.2 28.3 47.6 25.5 46.8 23.7 48.0 21.5 20.9 0.1 17.4 0.5 17.3 1.4 14.1
GraphPrompter 63.2 37.3 65.3 34.5 69.5 51.6 69.7 51.0 37.9 2.2 27.4 2.3 27.3 1.0 7.2
GraphGPT 62.4 39.6 65.0 41.7 71.2 61.6 62.2 38.9 43.2 16.2 25.4 1.7 24.3 2.0 6.8
LLaGA 62.0 39.6 52.2 28.0 49.4 30.4 48.8 18.0 28.0 0.2 18.2 0.9 16.6 1.6 12.4
SimGCL (Ours) 78.0 67.6 78.0 63.8 68.8 64.1 81.2 71.3 69.7 62.7 31.8 10.3 36.3 6.8 2.2
Obs. ❶ GNN-based methods exhibit persistent limitations in both NCIL and FSNCIL scenarios. As shown in Table 1 and Table 2, GNN-based approaches consistently underperform across all seven datasets. Notably, even on the small-scale Cora and Citeseer datasets, their final incremental accuracy ($\mathcal{A}_{N}$) fails to exceed 50%. This challenge even escalates significantly in larger-scale datasets like Arxiv-23 and Arxiv, where GCN and LwF achieve merely 1.5% and 1.6%, 2.0% and 3.8% accuracy after the final session of NCIL and FSNCIL, respectively, underscoring severe catastrophic forgetting. Moreover, classic CL methods (e.g., EWC and LwF) demonstrate inferior performance to vanilla GCN on shorter-session datasets and yield only marginal improvements in longer-session scenarios, while TPP fails to accurately identify task IDs, resulting in severely degraded performance under global testing settings. These limitations underscore the critical need for novel approaches to address GCIL and GFSCIL challenges effectively.
Obs. ❷ LLMs demonstrate higher performance in both NCIL and FSNCIL tasks, even without explicit graph structure utilization. As shown in Table 1 and Table 2, while the encoder-only BERT and RoBERTa achieve performance only comparable to or marginally better than GCN, the decoder-only LLaMA attains a 16.9% improvement on Citeseer in the FSNCIL scenario, while SimpleCIL surpasses all the GNN-based models in most datasets and metrics (24 out of 28), with average improvements of 12.93% in NCIL and 9.7% in FSNCIL, respectively. We attribute this superior performance to the LLM backbone’s exceptional generalization capability for processing textual features and capturing class-specific semantics. The results also demonstrate the substantial potential of LLMs in addressing NCIL and FSNCIL challenges.
[…]
Obs. ❹ Dense graph structures may enhance GLM effectiveness in GCL. Table 1 and Table 2 reveal that LLM-as-Predictor GLMs achieve notable performance gains on the WikiCS and Photo datasets, with GraphPrompter and GraphGPT ranking as runner-up or third-best models. This improvement likely stems from these datasets’ higher graph density than datasets like Cora or Products, providing richer structural information for LLM to better understand the task scenarios and make predictions. However, all GLMs exhibit suboptimal performance on Arxiv despite their comparable edge density (6.89 edges/node), suggesting that extended session ranges may disproportionately compromise LLMs’ generalization capacity, outweighing potential structural benefits.
Obs. ❺ Prototype-based learning improves cross-task generalization for GNNs and LLMs. Table 1 and Table 2 demonstrate that prototype-based methods achieve superior performance, with Cosine and SimpleCIL emerging as optimal approaches for GNN- and LLM-based methods, respectively. In NCIL settings with abundant per-session samples, these methods outperform other baselines by significant margins (maximum improvements of 29.7% and 39.2% versus GCN and RoBERTa). This advantage stems from an optimal balance between generalization and plasticity: Initial session tuning bridges the distribution gap between the initialized or pretrained representations and target tasks, while subsequent session prototype generation preserves model parameters, maintaining generalization without catastrophic interference. Moreover, the prototype framework requires only first-session training, offering significant computational efficiency benefits.
Figure 4: The performance of SimpleCIL and SimGCL across varying LLM backbone sizes on Arxiv. The ’B’ and ’Ro’ refer to BERT and RoBERTa.
Obs. ❻ SimGCL consistently overperform current GNN-, LLM- and GLM-based baselines. From Table 1 and Table 2, it is obvious that SimGCL consistently overperform other baselines (23 out of 28), with a maximum 21.7% and 18.0% improvement in NCIL and FSNCIL, respectively. The effectiveness of SimGCL stems from two key factors: (1) its graph-structured instruction tuning and prompting framework enhances LLMs’ comprehension of graph topology, yielding higher-quality prototypes enriched with task-specific structural information for more discriminative representations; (2) unlike GraphPrompter and GraphGPT that rely on GNN-derived representations, SimGCL encodes graph information through textual prompts that better align with LLMs’ native input space, eliminating cross-architecture representation mismatches. However, SimGCL demonstrates relatively inferior performance on the arxiv-23 dataset and in FSNCIL compared to NCIL, which we attribute to two factors: (1) The sparse graph structure of Arxiv-23 provides limited topological information, potentially compromising LLMs’ structural comprehension zhu2024efficient ; (2) SimGCL’s expanded tuning set (12 classes vs. 4 classes in NCIL) in FSNCIL may promote overfitting to the initial session data, thereby diminishing generalization capability for subsequent incremental sessions.
Obs. ❼ Scaling LLM parameters enhances generalization. Figure 4 compares the performance of SimpleCIL and SimGCL across LLM backbones of varying parameter scales. The results indicate that increased parameter counts consistently enhance generalization capacity across both encoder-only and decoder-only architectures, consequently boosting GCL performance throughout all task sessions.
Obs. ❽ LLMs and GLMs are better at dealing with tasks with short-range sessions. Table 3 evaluates model performance on the Arxiv dataset across varying session and class configurations [^2] Detailed results for BERT, GraphGPT, and LLaGA are provided in Appendix D.. In the 8-way 5-session setting, most LLM- and GLM-based methods achieve performance comparable to or exceeding top GNN-based approaches. However, this advantage diminishes with increasing session numbers, where performance differences between LLM/GLM and GNN methods become negligible. We attribute this convergence to progressive overfitting during incremental sessions, which increasingly weakens the pretrained models’ generalization capacity, ultimately reducing their performance advantage. Furthermore, prototype-based methods (i.e., Cosine, SimpleCIL, and SimGCL) demonstrate consistent performance stability across all experimental configurations, as their training-free prototype generation mechanism remains unaffected by session variations. This property makes them particularly suitable for long-session GCL applications.
Table 3: Performance comparison of LLM4GCL baselines in NCIL scenario across different class sizes (W) and session numbers (S) on Arxiv. The best, second, and third results are highlighted.
Methods 8W5S 5W8S 4W10S 2W20S Avg.
$\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ Rank
GCN 24.5 2.0 20.5 1.1 21.2 1.5 13.4 0.7 9.6
EWC 31.6 19.6 26.2 13.8 24.9 13.8 15.4 4.2 5.9
LwF 29.2 8.6 22.6 2.7 18.9 1.6 14.2 1.0 7.6
Cosine 35.6 23.2 38.3 27.9 37.4 27.8 41.6 31.4 2.9
RoBERTa 33.9 3.9 26.7 2.1 24.4 1.4 15.6 0.3 7.3
LLaMA 35.0 3.8 27.6 1.5 24.9 1.4 15.9 0.3 6.8
SimpleCIL 46.1 31.4 49.7 35.9 50.6 36.5 52.6 39.1 1.5
[…]
ENGINE 34.1 3.7 26.5 2.0 24.6 1.3 15.7 0.2 8.3
GraphPrompter 35.1 2.9 27.8 2.0 24.8 1.4 16.8 0.4 6.4
SimGCL (Ours) 51.6 28.7 53.0 30.4 59.9 33.8 57.4 17.5 1.6
5 Related Works
Continual Learning with Pre-trained Models (PTMs). The application of PTMs to continual learning was initially explored in Computer Vision (CV) zhou2025revisiting ; zhou2024continual and Natural Language Processing (NLP) wang2023orthogonal ; zhao2024sapt , with Vision Transformer (ViT) dosovitskiy2020image and BERT devlin2019bert serving as representative implementations in each domain respectively. Pre-trained on extensive corpora brown2020language rather than trained from scratch, PTMs exhibit inherent advantages in generalizability cao2023retentive , which supports flexible algorithm design, including both frozen-backbone approaches in a non-continual manner (e.g., training-free prototypes zhou2025revisiting ) and parameter-efficient tuning (PEFT) methods that preserve pre-trained weights (e.g., prompt tuning jia2022visual and adapter learning chen2022adaptformer ). Furthermore, leveraging the lightweight nature of PEFT blocks, there are also methods that employ orthogonal optimization via auxiliary losses wang2023orthogonal or null-space projection lu2024visual to minimize cross-task interference.
Graph Continual Learning (GCL). While continual learning in CV and NLP focuses on preventing catastrophic forgetting of semantic representations, GCL must additionally address the forgetting of learned topological patterns induced by evolving graph structures zhang2024continual . Typically, GCL methods can be categorized based on their approach to mitigating catastrophic forgetting, including (1) regularizing parameter updates through explicit constraints, such as topology weight liu2021overcoming or knowledge distillation hoang2023universal ; li2017learning ; (2) isolating task-critical parameters, typically implemented through expanding parameters or model components to accommodate new knowledge zhang2023continual ; zhang2022hierarchical ; and (3) replaying representative historical data, including key nodes zhou2021overcoming , sparsified computational subgraphs zhang2022sparsified ; zhang2023ricci , and condensed task-specific graphs liu2023cat ; niu2024graph . However, practical constraints like privacy concerns and storage limitations often prevent access to historical task data, thereby restricting the applicability of replay-based methods. Therefore, our benchmark primarily focuses on replay-free approaches.
Graph-enhanced Language Models (GLMs). Recently, with the rapid development of LLMs (e.g., RoBERTa conneau2019unsupervised , DeBERTa he2020deberta , LLaMA touvron2023llama , GPTs achiam2023gpt ), research interest in leveraging LLMs for graph tasks has grown substantially, resulting in numerous GLM approaches. Generally, these methods can be categorized into three main approaches based on how LLMs interact with graph structures or GNNs jin2024large ; li2023survey : (1) LLM-as-Enhancer he2023harnessing ; liu2023one ; zhu2024efficient , which leverages LLMs to retrieve node-relevant semantic knowledge, enhancing GNNs’ initial embedding quality; (2) LLM-as-Predictor chen2024llaga ; huang2024can ; li2024zerog , which employs LLMs to autoregressively make predictions using graph-structured information, including structural prompts and GNN-encoded representations; and (3) GNN-LLM Alignment jin2023patton ; zhao2022learning , which aligns LLMs with GNNs in the same vector space, enabling GNNs to be more semantically aware. Considering the generalization and semantic understanding capabilities of LLMs, combined with the graph structure modeling capacity of GNNs and other integrations, GLMs are anticipated to provide a novel approach to more effectively addressing GCL tasks.
6 Conclusion
This paper introduces LLM4GCL, the first comprehensive benchmark for evaluating GCL performance of LLMs and graph-enhanced LLMs. Our work makes three key contributions: Firstly, we systematically analyze the current GCL evaluation paradigm, identifying its critical flaws that may cause task ID leakage and compromise performance assessment. (2) We evaluate 9 representative LLM-based methods across 7 textual-attributed graphs, revealing that existing GLM-based approaches underperform in both NCIL and FSNCIL scenarios due to overfitting tendencies and LLM-GNN representation misalignment. However, prototype-based modifications enable these models to achieve state-of-the-art performance, with additional gains attainable through model scaling. (3) We propose SimGCL, an effective GLM-based approach combining graph-prompted instruction tuning with prototype networks, which enhances topological understanding while preserving LLMs’ generalization capacity. We believe the proposal of LLM4GCL will establish a foundation for future research by delineating effective pathways for PTM-graph integration. Moreover, we have made our code publicly available and welcome further contributions of new datasets and methods.
Appendix A Additional Details on LLM4GCL
A.1 Discussion on Local Testing and Global Testing
In this section, we will first introduce the pipeline of TPP [42], followed by a comprehensive analysis of the two evaluation paradigms: local testing and global testing, supported by empirical results.
A.1.1 Approaches
TPP [^1] In this work, we focus solely on formalizing the pipeline of TPP; for rigorous mathematical derivations and proofs, we direct readers to the original literature.. TPP follows the local testing manner and proposes to use a Laplacian smoothing approach to generate a prototypical embedding for each graph task for task ID prediction. As for the $i$-th task with graph data $\mathcal{G}_{s_{i}}=\mathcal{G}_{q_{i}}=(\mathbf{A}_{s_{i}},\mathbf{X}_{s_{i}},\mathcal{V}_{s_{i}}^{l},\mathcal{V}_{s_{i}}^{u})$, TPP constructed a task prototype $\mathbf{p}_{s_{i}}$ based on the train set $\{\mathbf{x}_{j}\mid v_{j}\in\mathcal{V}_{s_{i}}^{l}\}$. Specifically, given training graph $\mathcal{G}_{s_{i}}$, the Laplacian smoothing is first applied to obtain the node embeddings $\mathbf{Z}_{s_{i}}$:
$\mathbf{Z}_{s_{i}}=(\mathbf{I}-(\mathbf{\hat{D}}_{s_{i}})^{-\frac{1}{2}}\mathbf{\hat{L}}_{s_{i}}(\mathbf{\hat{D}}_{s_{i}})^{-\frac{1}{2}})^{k}\mathbf{X}_{s_{i}}$ (4)
where $k$ is the smoothing steps, $\mathbf{I}$ is an identity matrix, $\mathbf{\hat{L}}_{s_{i}}$ is the graph Laplacian matrix of $\mathbf{\hat{A}}_{s_{i}}$, and $\mathbf{\hat{D}}_{s_{i}}$ is the diagonal degree matrix. Consequently, the training task prototype $\mathbf{p}_{s_{i}}$ can be generated as:
$\mathbf{p}_{s_{i}}=\frac{1}{|\mathcal{V}_{s_{i}}^{l}|}\sum_{v_{j}\in\mathcal{V}_{s_{i}}^{l}}\mathbf{z}_{s_{i},j}(\mathbf{\hat{D}}_{s_{i},jj})^{-\frac{1}{2}}$ (5)
Thus, all the task prototypes can be separately constructed and stored as $\mathcal{P}=\{\mathbf{p}_{s_{1}},\mathbf{p}_{s_{2}},...,\mathbf{p}_{s_{n}}\}$. Similarly, the task prototype of the local testing on $\mathcal{G}_{q_{i}}=\mathcal{G}_{s_{i}}$ can be denoted as:
$\mathbf{p}_{q_{i}}=\frac{1}{|\mathcal{V}_{q_{i}}^{u}|}\sum_{v_{j}\in\mathcal{V}_{q_{i}}^{u}}\mathbf{z}_{q_{i},j}(\mathbf{\hat{D}}_{q_{i},jj})^{-\frac{1}{2}}$ (6)
Then, during local testing, the task ID prediction is conducted between the train and test prototypes:
$t_{q_{i}}=\arg\min(d(\mathbf{p}_{q_{i}},\mathbf{p}_{s_{1}}),d(\mathbf{p}_{q_{i}},\mathbf{p}_{s_{2}}),...,d(\mathbf{p}_{q_{i}},\mathbf{p}_{s_{n}}))$ (7)
After precisely identifying the task ID, TPP subsequently incorporates a graph prompt learning framework that assigns distinct prompt vectors to adapt a pretrained GNN for diverse tasks:
$\overline{\mathbf{x}}_{q_{i},j}=\mathbf{x}_{q_{i},j}+\sum_{m}^{M}\alpha_{m}\phi_{i,m},\alpha_{m}=\frac{e^{(\mathbf{w}_{m})^{T}\mathbf{x}_{q_{i},j}}}{\sum_{l}^{M}e^{(\mathbf{w}_{l})^{T}\mathbf{x}_{q_{i},j}}}$ (8)
where $\Phi_{i}=[\phi_{i,1},\phi_{i,2},...,\phi_{i,M}]^{T}\in\mathbb{R}^{M\times F}$ represents the learnable graph prompt for the $i$-th task, with $M$ denoting the number of vector-based tokens $\phi_{i}$. $\alpha_{m}$ is the importance score of token $\phi_{i,m}$ in the prompt, while $\mathbf{w}_{j}$ corresponds to a learnable projection weight. The final prompted graph representation $\mathcal{G}_{q_{i}}=\{\mathbf{A}_{q_{i}},\mathbf{X}_{q_{i}}+\Phi_{i}\}$ is then processed by a frozen pre-trained GNN model $f(\cdot,\cdot)$ to generate predictions:
$\hat{\mathcal{Y}}_{q_{i}}=\varphi_{i}(f(\mathbf{A}_{q_{i}},\mathbf{X}_{q_{i}}+\Phi_{i}))$ (9)
where $\varphi_{i}$ denotes the task-specific MLP layer for the $i$-th task. Crucially, TPP effectively transforms the class-incremental learning problem into a task-incremental learning paradigm, decomposing the overall continual learning objective into multiple independent sub-tasks, thereby significantly mitigating the learning difficulty.
Mean Pooling. TPP proposes that when applying a sufficiently large number of Laplacian smoothing iterations $k$, the distance between training and test prototypes from the same task approaches zero, while maintaining substantial separation between prototypes belonging to different CIL tasks. This discriminative property ensures reliable prediction accuracy. To evaluate the effectiveness of this technique, we implement an ablation model called Mean Pooling, where the Laplacian smoothing operation is replaced by a basic mean pooling aggregation, formally defined as:
$\mathbf{p}_{s_{i}}=\frac{1}{|\mathcal{V}_{s_{i}}^{l}|}\sum_{v_{j}\in\mathcal{V}_{s_{i}}^{l}}\mathbf{x}_{s_{i},j},\ \ \mathbf{p}_{q_{i}}=\frac{1}{|\mathcal{V}_{q_{i}}^{u}|}\sum_{v_{j}\in\mathcal{V}_{q_{i}}^{u}}\mathbf{x}_{q_{i},j}$ (10)
Under this formulation, the distance between training and test prototypes depends solely on the feature distributions of the training and testing nodes.
MLP. By decomposing the CIL task into independent subtasks, TPP eliminates the need for session-wide classification across all accumulated classes. Instead, the model can focus exclusively on individual subtasks, significantly reducing task complexity. Building upon this observation, we propose an additional ablation model employing a basic two-layer MLP for classification:
$\hat{\mathcal{Y}}_{q_{i}}=\varphi_{i}(\mathbf{X}_{q_{i}})$ (11)
A.1.2 Experimental Results
Table 4: Performance evaluation of TPP against two ablation models: Mean Pooling and MLP. AA and AF are the average classification accuracy and forgetting ratio across all sessions, respectively.
Methods Cora Citeseer WikiCS Photo Products Arxiv-23 Arxiv
AA AF AA AF AA AF AA AF AA AF AA AF AA AF
TPP 95.2 -0.0 87.9 -0.0 93.6 -0.0 91.3 -0.0 89.6 -0.0 91.6 -0.0 85.7 -0.0
Mean Pooling 95.2 -0.0 87.9 -0.0 93.6 -0.0 91.3 -0.0 89.6 -0.0 91.6 -0.0 85.7 -0.0
MLP 90.3 -0.0 86.9 -0.0 89.2 -0.0 79.0 -0.0 83.1 -0.0 88.8 -0.0 81.6 -0.0
To investigate the impact of Laplacian smoothing and assess subtask complexity, we evaluate the performance of TPP against two variants, Mean Pooling and MLP, with results presented in Table 4. We employ the local testing paradigm, where model performance is quantified using two key metrics: Average Accuracy (AA) and Average Forgetting (AF). These metrics are defined as follows:
$\text{AA}=\frac{1}{n}\sum_{j=1}^{n}(\mathcal{A}_{n,j}),\ \ \text{AF}=\frac{1}{n}\sum_{j=1}^{n-1}(\mathcal{A}_{n,j}-\mathcal{A}_{j,j})$ (12)
where $n$ denotes the total number of sessions, and $\mathcal{A}_{i,j}$ represents the model’s classification accuracy for task $\mathcal{T}_{q_{j}}$ after training on task $\mathcal{T}_{s_{i}}$. The experimental results yield two key conclusions:
♣ Laplacian smoothing demonstrates limited necessity for accurate task identification. Comparative analysis between TPP and Mean Pooling demonstrates that both approaches achieve accurate task ID prediction. This phenomenon can be explained by the local testing setup, where training and testing nodes are drawn from the same graph ($\mathcal{G}_{s_{i}}=\mathcal{G}_{q_{i}}$). In this configuration, the feature distribution similarity between training and testing nodes within each class provides sufficient discriminative information for reliable task identification. These results indicate that the local testing paradigm is inherently susceptible to task ID leakage.
♠ Decomposition into subtasks significantly reduces task complexity. As evidenced by the experimental results, the MLP baseline achieves comparable performance to TPP across Citeseer, WikiCS, Arxiv-23, and Arxiv datasets, with merely a 3.08% average performance gap. This finding demonstrates that task decomposition effectively reduces learning difficulty, enabling simpler architectures to attain competitive performance. However, this characteristic may compromise the evaluation of models’ true continual learning capabilities, as the simplified subtasks fail to fully capture the challenges inherent in the original CIL problem.
A.2 Details of Datasets
Table 5: Statistics and experimental settings for all datasets in LLM4GCL. The symbol ’’ indicates adjusted statistics that differ from the original dataset.
Dataset Cora Citeseer WikiCS Photo Products Arxiv-23 Arxiv
Attributes # Nodes 2,708 3,186 11,701 48,362 53,994 46,196 169,343
# Edges 5,429 4,277 215,863 500,928 72,242 78,542 1,166,243
# Features 1,433 3,703 300 768 100 300 128
# Classes 7 6 10 12 31 37 40
Avg. # Token 183.4 210.0 629.9 201.5 152.9 237.7 239.8
Domain Citation Citation Web link E-Commerce E-Commerce Citation Citation
NCIL # Classes per session 2 2 3 3 4 4 4
# Sessions 3 3 3 4 8 9 10
# Shots per class 100 100 200 400 400 400 800
FSNCIL # Base Classes 3 2 4 4 11 13 12
# Novel Classes 4 4 6 8 20 24 28
# Ways per session 2 2 3 4 4 4 4
# Sessions 3 3 3 3 6 7 8
# Shots per base class 100 100 200 400 400 400 800
# Shots per novel class 5 5 5 5 5 5 5
All of the public datasets used in LLM4GCL were previously published, covering a multitude of domains. For each dataset, we store the graph-type data in the .pt format using PyTorch. This includes shallow embeddings, raw text of nodes, edge indices, node labels, and label names. The statistics of these datasets are presented in Table 5, and the detailed descriptions are listed in the following:
  • Cora [8]. The Cora dataset is a citation network comprising research papers and their citation relationships within the computer science domain. The raw text data for the Cora dataset was sourced from the GitHub repository provided in Chen et al. [^2] https://github.com/CurryTang/Graph-LLM. In this dataset, each node represents a research paper, and the raw text feature associated with each node includes the title and abstract of the respective paper. An edge in the Cora dataset signifies a citation relationship between two papers. The label assigned to each node corresponds to the category of the paper.
  • Citeseer [8]. The Citeseer dataset is a citation network comprising research papers and their citation relationships within the computer science domain. The TAG version of the dataset contains text attributes for 3,186 nodes, and the raw text data for the Citeseer dataset was sourced from the GitHub repository provided in Chen et al.. Each node represents a research paper, and each edge signifies a citation relationship between two papers.
  • WikiCS [32]. The WikiCS dataset is an internet link network where each node represents a Wikipedia page, and each edge represents a reference link between pages. The raw text data for the WikiCS dataset was collected from OFA [^3] https://github.com/LechengKong/OneForAll. The raw text associated with each node includes the name and content of a Wikipedia entry. Each node’s label corresponds to the category of the entry.
  • Ele-Photo (Photo) [55]. The Ele-Photo dataset [^4] https://github.com/sktsherlock/TAG-Benchmark is derived from the Amazon Electronics dataset [40]. The nodes in the dataset represent electronics-related products, and edges between two products indicate frequent co-purchases or co-views. The label for each dataset corresponds to the three-level label of the electronics products. User reviews on the item serve as its text attribute. In cases where items have multiple reviews, the review with the highest number of votes is utilized. For items lacking highly-voted reviews, a user review is randomly chosen as the text attribute.
  • Ogbn-products (Products) [15]. The OGBN-Products dataset is characterized by its large scale, originally containing approximately 2 million nodes and 61 million edges. Following the node sampling strategy proposed in TAPE [^5] https://github.com/XiaoxinHe/TAPE, we utilize a subset consisting of about 54k nodes and 72k edges for experimental use. In this dataset, each node represents a product sold on Amazon, and edges between two products indicate frequent co-purchasing relationships. The raw text data for each node includes product titles, descriptions, and/or reviews.
  • Arxiv-23 [15]. The Arxiv-23 dataset is proposed in TAPE, representing a directed citation network comprising computer science arXiv papers published in 2023 or later. Similar to Ogbn-arxiv, each node in this dataset corresponds to an arXiv paper, and directed edges indicate citation relationships between papers. The raw text data for each node includes the title and abstract of the respective paper. The label assigned to each node represents one of the 40 subject areas of arXiv CS papers, such as cs.AI, cs.LG, and cs.OS. These subject areas are manually annotated by the paper’s authors and arXiv moderators.
  • Ogbn-arxiv (Arxiv) [32]. The Ogbn-arxiv dataset is a citation network comprising papers and their citation relationships, collected from the arXiv platform. The raw text data for the dataset was sourced from the GitHub repository provided in OFA. The original raw texts are available here [^6] https://snap.stanford.edu/ogb/data/misc/ogbn_arxiv. Each node represents a research paper, and each edge denotes a citation relationship.
A.3 Details of Baseline Models
In LLM4GCL, we integrate 15 baseline methods for NCIL and FSNCIL tasks, including 6 GNN-based methods, 4 LLM-based methods, and 5 GLM-based methods. We provide detailed descriptions of these methods used in our benchmark as follows.
  • GNN-based methods.
    • GCN [24]. The first deep learning model that leverages graph convolutional layers. In LLM4GCL, we employ a two-layer GCN followed by one MLP layer. We use the code available at https://github.com/pyg-team/pytorch_geometric.
    • EWC [25]. A regularization-based CL method that prevents catastrophic forgetting through quadratic penalties on key parameter deviations. Parameter importance is quantified via the Fisher information matrix, maintaining performance on learned tasks during new task acquisition. We use the code available at https://github.com/QueuQ/CGLB.
    • LwF [31]. LwF minimizes the discrepancy between the logits of the old model and the new model through knowledge distillation to preserve knowledge from the old tasks. We use the code available at https://github.com/QueuQ/CGLB.
    • Cosine. Cosine first trains a GNN on the initial task session, subsequently freezes the network parameters, and leverages a training-free prototype mechanism to generate task-specific prototypes, ultimately employing cosine similarity for classification.
    • TPP [42]. A replay-and-forget-free GCL approach that employs Laplacian smoothing-based task identification to precisely predict task IDs and incorporates graph prompt engineering to adaptively transform the GNN into a series of task-specific classification modules. We use the code available at https://github.com/mala-lab/TPP.
    • TEEN [49]. TEEN is a training-free prototype calibration strategy that addresses the issue of new classes being misclassified into base classes by leveraging the semantic similarity between base and new classes, which is captured by a feature extractor pre-trained on base classes. We use the code available at https://github.com/wangkiw/TEEN.
  • LLM-based methods.
    • BERT [2, 10]. A Transformer-based model family that employs masked language modeling and next sentence prediction pretraining to learn bidirectional contextual representations. It achieves state-of-the-art performance across multiple NLP tasks with minimal task-specific modifications. We use the code available at https://github.com/google-research/bert and https://github.com/prajjwal1/generalize_lm_nli.
    • RoBERTa [35]. An optimized BERT variant family that enhances pretraining through dynamic masking, larger batches, and extended data. It removes the next sentence prediction objective while maintaining MLM, achieving superior performance across NLP benchmarks with more efficient training. We use the code available at https://github.com/facebookresearch/fairseq/tree/main/examples/xlmr.
    • LLaMA [48]. A foundational large language model family featuring architectural optimizations like RMSNorm and SwiGLU activations. Trained on trillions of tokens from diverse public corpora, it achieves remarkable few-shot performance while maintaining efficient inference compared to similarly-sized models. We use the code available at https://github.com/meta-llama/llama-cookbook.
    • SimpleCIL [67]. A computationally efficient approach that freezes the PTM’s embedding function and derives class prototypes by averaging embeddings for classification, which harnesses the model’s inherent generalizability for effective knowledge transfer without fine-tuning. We use the code available at https://github.com/LAMDA-CL/RevisitingCIL.
  • GLM-based methods.
[…]
    • ENGINE [70]. A framework adds a lightweight and tunable G-Ladder module to each layer of the LLM, which uses a message-passing mechanism to integrate structural information. This enables the output of each LLM layer (i.e., token-level representations) to be passed to the corresponding G-Ladder, where the node representations are enhanced and then used for node classification. We use the code available at https://github.com/ZhuYun97/ENGINE.
    • GraphPropmter [36]. A framework integrating graph structures and LLMs via adaptive prompts, projecting GNN-based structural embeddings through MLP to align with the LLM’s semantic space for joint topology-text processing. We use the code available at https://github.com/franciscoliu/graphprompter.
    • GraphGPT [46]. A framework that initially aligns the graph encoder with natural language semantics through text-graph grounding, and then combines the trained graph encoder with the LLM using a projector, through which the model can directly complete graph tasks with natural language, thus performing zero-shot transferability. We use the code available at https://github.com/WxxShirley/LLMNodeBed.
    • LLaGA [6]. LLaGA uses node-level templates to transform graph data into structured sequences, then maps them into token embeddings, enabling LLMs to process graph data with improved generalizability. We use the code available at https://github.com/WxxShirley/LLMNodeBed.
Appendix B Details on Implementations
B.1 Hyper-parameters
  • For GCN, we grid-search the hyperparameters
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256]
        layer_num = 2, dropout = 0.5
  • For EWC, we grid-search the hyperparameters
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256],
        strength in [1, 100, 10000]
        layer_num = 2, dropout = 0.5
  • For LwF, we grid-search the hyperparameters
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256],
        strength in [1, 100, 10000], lambda in [0.1, 1.0], T in [0.2, 2.0]
        layer_num = 2, dropout = 0.5
  • For Cosine, we grid-search the hyperparameters
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256]
        layer_num = 2, dropout = 0.5, T = 1.0, sample_num = 100
  • For TPP, we grid-search the hyperparameters
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256],
        pe in [0.2, 0.3], pf in [0.2, 0.3]
        layer_num = 2, dropout = 0.5,
        pretrain_batch_size = 500, pretrain_lr = 1e-3
  • For TEEN, we grid-search the hyperparameters
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256]
        layer_num = 2, dropout = 0.5
        T = 1.0, sample_num = 100, softmax_T = 16, shift_weight = 0.5
  • For BERT and RoBERTa, we grid-search the hyperparameters
        lr in [1e-4, 2e-4, 5e-4], batch_size in [5, 10, 20]
        min_lr = 5e-6, weight_decay = 5e-2,
        dropout = 0.1, att_drouput = 0.1, max_length = 256
        lora_r = 5, lora_alpha = 16, lora_dropout = 0.05
  • For LLaMA, we grid-search the hyperparameters
        lr in [1e-4, 2e-4, 5e-4], batch_size in [5, 10, 20]
        min_lr = 5e-6, weight_decay = 5e-2,
        dropout = 0.1, att_drouput = 0.1, max_length = 512
        lora_r = 5, lora_alpha = 16, lora_dropout = 0.05
  • For SimpleCIL, we grid-search the hyperparameters
        lr in [1e-4, 2e-4, 5e-4], batch_size in [5, 10, 20]
        min_lr = 5e-6, weight_decay = 5e-2,
        dropout = 0.1, att_drouput = 0.1, max_length = 256
        lora_r = 5, lora_alpha = 16, lora_dropout = 0.05
        T = 1.0, sample_num = 20
[…]
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256]
        layer_num = 2, dropout = 0.5
  • For ENGINE, we grid-search the hyperparameters
        lr in [1e-5, 1e-4, 1e-3], hidden_dim in [64, 128, 256], r in [1, 32]
        layer_num = 1, dropout = 0.5, T = 0.1,
        layer_select = [0, 6, 12, 18, 24, -1]
  • For GraphPrompter, we follow the instructions from the paper [36], use
        lr = 2e-4, batch_size = 10, min_lr = 5e-6, weight_decay = 5e-2,
        dropout = 0.1, att_drouput = 0.1, max_length = 512
        lora_r = 5, lora_alpha = 16, lora_dropout = 0.05
        gnn = ’GCN’, layer_num = 4, hidden_dim = 1024, output_dim = 1024,
        dropout = 0.5, proj_hidden_dim = 1024
  • For GraphGPT, we follow the instructions from the paper [54], use
        dropout = 0.1, att_drouput = 0.1, max_length = 512
        s1_k_hop = 2, s1_num_neighbors = 5, s1_max_txt_length = 512,
        s1_epoch = 2, s1_batch_size = 5, s1_lr = 1e-4
        s2_num_neighbors = 4, max_txt_length = 512,
        s2_epoch = 10, s2_batch_size = 5, s2_lr = 1e-4
        lora_r = 5, lora_alpha = 16, lora_dropout = 0.05
  • For LLaGA, we follow the instructions from the paper [6], use
        lr = 2e-4, batch_size = 10, min_lr = 5e-6, weight_decay = 5e-2,
        dropout = 0.1, att_drouput = 0.1, max_length = 512
        llm_freeze = ’True’, neighbor_template = ’ND’, nd_mean = True,
        k_hop = 2, sample_size = 10, hop_field = 4, proj_layer = 2
  • For SimGCL, we follow the instructions from the paper [51], use
        lr = 2e-4, batch_size = 10, min_lr = 5e-6, weight_decay = 5e-2,
        lora_r = 5, lora_alpha = 16, lora_dropout = 0.05,
        T = 1.0, sample_num = 50, hop = [20, 20],
        include_label = False, max_node_text_len = 128
Table 6: Selection of GNN and LLM Backbones in LLM4GCL, $\mathbf{A},\mathbf{X},\mathcal{R}$ refers to the adjacent matrix, feature matrix, and raw text attributes of the input TAG, respectively.
Type Method Predictor Input Output Format GNN Backbone LLM Backbone
GNN-based GCN [24] GNN $(\mathbf{A,\mathbf{X}})$ Logits GCN -
EWC [25] GNN $(\mathbf{A,\mathbf{X}})$ Logits GCN -
LwF [31] GNN $(\mathbf{A,\mathbf{X}})$ Logits GCN -
Cosine GNN $(\mathbf{A,\mathbf{X}})$ Logits GCN -
TPP [42] GNN $(\mathbf{A,\mathbf{X}})$ Logits SGC, GCN -
TEEN [49] GNN $(\mathbf{A,\mathbf{X}})$ Logits GCN -
LLM-based BERT [10] LLM $\mathcal{R}$ Logits - BERT-base (110M)
RoBERTa [35] LLM $\mathcal{R}$ Logits - RoBERTa-large (355M)
LLaMA [48] LLM $\mathcal{R}$ Prediction Texts - LLaMA3-8B
SimpleCIL [67] LLM $\mathcal{R}$ Logits - RoBERTa-large (355M)
[…]
ENGINE [70] GNN $(\mathbf{A},\mathcal{R})$ Logits GCN LLaMA3-8B
GraphPrompter [20] LLM $(\mathbf{A},\mathbf{X},\mathcal{R})$ Prediction Texts GCN LLaMA3-8B
GraphGPT [46] LLM $(\mathbf{A},\mathbf{X},\mathcal{R})$ Prediction Texts - LLaMA3-8B
LLaGA [6] LLM $(\mathbf{A},\mathbf{X},\mathcal{R})$ Prediction Texts - LLaMA3-8B
B.2 Backbone Selection
To compare the various baseline methods in LLM4GCL as fairly as possible, we try to utilize the same GNN and LLM backbones in our implementations. Table 6 shows the GNN and LLM backbones used in the original implementations, as well as those implemented in LLM4GCL.
Appendix C Text Prompt Design
[…]
Appendix D Additional Experimental Results
D.1 Session Numbers
Table 7: Performance comparison of LLM4GCL baselines in NCIL scenario across different class sizes (W) and session numbers (S) on Arxiv. The best, second, and third results are highlighted.
Methods 8W5S 5W8S 4W10S 2W20S Avg.
$\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ Rank
GCN 24.5 2.0 20.5 1.1 21.2 1.5 13.4 0.7 12.5
EWC 31.6 19.6 26.2 13.8 24.9 13.8 15.4 4.2 7.6
LwF 29.2 8.6 22.6 2.7 18.9 1.6 14.2 1.0 10.2
Cosine 35.6 23.2 38.3 27.9 37.4 27.8 41.6 31.4 3.1
BERT 36.7 15.0 29.7 10.2 26.0 9.7 16.6 5.0 5.1
RoBERTa 33.9 3.9 26.7 2.1 24.4 1.4 15.6 0.3 10.1
LLaMA 35.0 3.8 27.6 1.5 24.9 1.4 15.9 0.3 9.6
SimpleCIL 46.1 31.4 49.7 35.9 50.6 36.5 52.6 39.1 1.5
[…]
ENGINE 34.1 3.7 26.5 2.0 24.6 1.3 15.7 0.2 11.2
GraphPrompter 35.1 2.9 27.8 2.0 24.8 1.4 16.8 0.4 8.9
GraphGPT 38.9 7.4 35.0 6.2 32.2 3.9 23.5 3.1 5.1
LLaGA 34.8 8.6 27.9 5.3 26.1 1.3 16.6 0.9 7.6
SimGCL (Ours) 51.6 28.7 53.0 30.4 59.9 33.8 57.4 17.5 1.6
Table 7 presents the performance of baseline models and SimGCL on the Arxiv dataset under different session-class configurations. Consistent with Obs. ❽, as session numbers increase, the performance advantages of BERT, GraphGPT, and LLaGA over GNN-based approaches diminish significantly, demonstrating pronounced catastrophic forgetting in long-session scenarios.
D.2 Extended Analysis of SimGCL
Table 8: Performance comparison of GNN-, LLM-, and GLM-based methods in the NCIL scenario of LLM4GCL. The best, second, and third results are highlighted.
Methods Cora Citeseer WikiCS Photo Products Arxiv-23 Arxiv Avg.
$\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ Rank
GCN 57.0 38.2 52.4 30.2 54.9 34.9 46.5 19.9 25.5 5.4 19.9 2.9 21.2 1.5 10.5
EWC 56.0 31.0 45.9 28.5 55.3 33.0 46.9 20.5 29.4 15.9 25.7 15.0 24.9 13.8 8.7
LwF 55.7 30.8 47.8 28.2 55.0 34.0 45.8 20.4 26.0 7.5 21.1 6.8 18.9 1.6 11.1
Cosine 65.4 45.2 50.7 31.0 66.5 53.5 63.6 49.6 36.1 16.1 36.1 24.2 37.4 27.8 4.3
TPP 45.7 13.7 45.4 9.6 35.1 9.8 36.3 5.7 15.0 0.0 12.2 0.3 19.6 3.4 15.3
BERT 56.0 29.9 53.8 28.7 58.4 30.0 43.4 18.4 26.9 4.1 27.1 5.1 26.0 9.7 9.9
RoBERTa 54.6 29.6 54.1 28.6 55.1 30.8 43.5 19.1 27.6 3.2 23.1 0.8 24.4 1.4 11.7
LLaMA 65.6 53.8 55.7 31.7 55.5 30.9 44.6 19.1 29.8 0.4 24.3 1.0 24.9 1.4 8.7
SimpleCIL 70.8 58.3 66.4 49.5 71.4 57.3 62.1 52.5 66.8 52.6 52.4 38.8 50.6 36.5 2.2
[…]
ENGINE 59.2 31.3 53.5 29.8 56.4 30.1 47.9 21.0 27.2 1.1 22.5 0.7 24.6 1.3 10.2
GraphPrompter 61.9 46.8 60.2 30.6 59.6 38.3 51.5 31.0 29.0 0.8 23.4 0.9 24.8 1.4 7.5
GraphGPT 55.5 31.6 60.0 30.1 62.0 49.2 50.8 30.2 35.5 3.2 30.7 5.8 32.2 3.9 6.6
LLaGA 58.2 30.2 51.3 27.8 53.7 27.6 47.2 20.7 25.7 0.2 26.9 4.1 26.1 1.3 11.3
SimGCL-1 Hops 89.4 83.9 76.4 64.2 32.4 20.3 44.7 33.5 70.9 43.7 28.2 6.1 38.2 10.7 5.5
SimGCL-2 Hops 84.6 80.0 77.1 66.3 73.5 61.9 82.1 72.6 71.1 60.2 38.7 13.6 59.9 33.8 1.5
Table 9: Performance comparison of GNN-, LLM-, and GLM-based methods in the FSNCIL scenario of LLM4GCL. The best, second, and third results are highlighted.
Methods Cora Citeseer WikiCS Photo Products Arxiv-23 Arxiv Avg.
$\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ $\mathcal{\bar{A}}$ $\mathcal{A}_{N}$ Rank
GCN 68.0 38.1 39.5 17.4 62.4 47.4 58.5 32.3 36.0 14.1 22.2 5.4 25.3 2.0 10.4
EWC 59.0 36.3 49.2 21.2 58.4 40.4 62.0 28.4 45.7 31.5 36.4 29.5 31.5 18.3 8.7
LwF 63.3 43.5 45.1 20.7 59.4 41.0 60.1 29.4 50.3 38.7 30.8 15.1 27.8 3.8 8.4
Cosine 72.6 57.8 49.1 25.7 68.0 50.7 67.9 50.5 50.9 33.6 40.1 27.9 27.2 17.2 5.2
TEEN 60.9 40.3 59.0 39.5 59.3 42.4 59.3 35.5 49.6 28.4 39.2 27.1 31.8 18.6 6.8
TPP 39.0 9.1 37.3 12.6 37.3 12.4 40.0 14.0 14.0 0.0 7.3 0.0 11.1 6.0 16.2
BERT 56.4 34.7 61.1 38.8 61.7 33.0 47.7 25.8 22.4 3.8 15.2 3.4 14.9 6.0 12.1
RoBERTa 59.6 41.9 54.0 29.2 67.2 42.3 58.2 29.6 38.8 6.9 22.6 1.4 25.7 1.3 10.4
LLaMA 72.6 55.6 75.9 55.5 65.2 43.9 61.4 32.3 43.5 13.7 23.5 1.5 22.7 1.4 7.8
SimpleCIL 69.6 53.6 64.1 49.9 73.2 63.1 66.3 53.0 65.6 53.6 49.8 40.0 46.4 36.6 2.6
[…]
ENGINE 52.2 28.3 47.6 25.5 46.8 23.7 48.0 21.5 20.9 0.1 17.4 0.5 17.3 1.4 15.0
GraphPrompter 63.2 37.3 65.3 34.5 69.5 51.6 69.7 51.0 37.9 2.2 27.4 2.3 27.3 1.0 7.8
GraphGPT 62.4 39.6 65.0 41.7 71.2 61.6 62.2 38.9 43.2 16.2 25.4 1.7 24.3 2.0 7.5
LLaGA 62.0 39.6 52.2 28.0 49.4 30.4 48.8 18.0 28.0 0.2 18.2 0.9 16.6 1.6 13.3
SimGCL-1 Hop 85.5 77.9 77.3 63.5 30.4 23.2 50.6 44.8 57.2 30.1 19.9 4.2 25.9 3.9 7.4
SimGCL-2 Hops 78.0 67.6 78.0 63.8 68.8 64.1 81.2 71.3 69.7 62.7 31.8 10.3 36.3 6.8 2.4
To comprehensively assess SimGCL’s performance, we conducted an evaluation using 1-hop neighborhood aggregation, with detailed results presented in Table 8 and Table 9.
Observation ❾ Rich structural information better enhances task comprehension. In Table 8 and Table 9, SimGCL exhibits performance degradation when limited to 1-hop neighborhood information, particularly on densely-connected datasets, WikiCS and Photo, where it underperforms even the vanilla GCN baseline. We hypothesize this stems from an inadequate structural context, misleading the model’s representation learning. Notably, modest improvements occur on sparser graph datasets, Cora and Citeseer, we attribute this to their smaller scale and lower connectivity makes 1-hop information sufficient for effective prediction.
D.3 Visualization
To have deeper insights into the comparative learning dynamics, we present performance evolution curves in Figure 5. The visualization demonstrates SimGCL’s consistent superiority over baseline models across multiple training sessions, underscoring its robust capability to maintain previously acquired knowledge while effectively assimilating new information.
Figure 5: Performance evolution of all models on Citeseer, Photo, Products, and Arxiv datasets.
Appendix E Broader Discussion
E.1 Limitations
Restricted to node-level tasks. Currently, LLM4GCL is limited to GCL for node classification tasks, as most GLM-based methods are designed for single-task scenarios, and node classification remains the predominant task in GML. However, GML encompasses other critical tasks such as link prediction and graph classification. Moreover, in practical applications, GCL has significant value in these task domains, like predicting interactions between existing and new users (edge-level class incremental learning, ECIL) and identifying novel protein categories (graph-level class incremental learning, GCIL). Notably, recent advances in Graph Foundation Models [39] have enabled Graph-enhanced LLMs like OFA [32], UniGraph [16], and UniAug [47] to support edge-level or graph-level tasks. Therefore, a comprehensive evaluation of GLM-based methods in ECIL and GCIL scenarios is imperative, and we designate this as one of the key directions for our future research.
Restricted to TAGs. In LLM4GCL, all seven benchmark datasets are TAGs, as the evaluated GLM-based methods inherently depend on textual node attributes. However, this requirement limits applicability to real-world non-textual graphs (e.g., transportation networks or flight routes). Currently, emerging solutions like GCOPE [64] and SAMGPT [57] demonstrate promising approaches for text-free graph foundation models. Consequently, extending GLM-based methods to such non-textual graph scenarios remains a critical direction for future research.
E.2 Potential Impacts
The integration of LLMs with graph tasks represents an emerging and highly promising research direction, demonstrating substantial potential across diverse applications and particularly in GCL scenarios. This paper proposes LLM4GCL to advance research focus on this novel paradigm that leverages graph-enhanced LLMs for improved GCL performance. Through LLM4GCL, we conduct a systematic investigation into graph-enhanced LLM methodologies and establish comprehensive benchmarking across various graph domains. We believe this work will serve as the catalyst and pioneer for accelerated progress in this developing research community.
Moreover, as LLM4GCL demonstrates several efficient LLM- and GLM-based GCL methods, we argue that these advancements hold significant societal value. Specifically, the proposed methods facilitate efficient knowledge updates in dynamic environments, such as social networks and recommendation systems, through seamless integration with LLMs, thereby substantially reducing computational overhead compared to conventional training-from-scratch paradigms. Furthermore, the cross-domain adaptability of these models offers distinct advantages in mission-critical applications, including healthcare diagnostics and financial forecasting. In such domains, continuous model adaptation to emergent knowledge, ranging from novel therapeutic interventions to evolving market trends, can be achieved while effectively mitigating catastrophic forgetting. However, practical deployment necessitates careful consideration of several key challenges, including data privacy implications, the potential for bias propagation in automated decision-making processes, and ensuring robustness against adversarial data injection, particularly in open-domain applications.
In the future, we will keep track of newly emerged techniques in the GCL field and continuously update LLM4GCL with more solid experimental results and detailed analyses.